OpenStudy (anonymous):

How many square inches of aluminum are required to make a soup can that has a diameter of 4 inches and a height of 8 inches? can you please show me the steps involved? thanks

OpenStudy (amistre64):

determine the surface area ...

OpenStudy (amistre64):

a soup can is made up of 3 surfaces: a bottom circle; a top circle; and a rectangle that is bent into a tube for the middle

OpenStudy (anonymous):

You need to find an area of cylinder: S=2*pi*r*(r+h) r=4/2=2, h=8 S=2*3.14*2(2+8)=125.6 sq.m.